OpenStudy (anonymous):

What would be the resulting molarity of a solution made by dissolving 45.7 grams of Mg(OH)2 in enough water to make a 1087-milliliter solution? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

@AccessDenied @mathstudent55 @Luigi0210 @Destinymasha @timo86m

OpenStudy (anonymous):

Well, we need to find the moles Mg(OH)2 45.7 g /58.319 g/mol=0.784 m = 0.784 / 1.087 L=0.721 I learned this yesterday, lol.